OpenStudy (anonymous):

express the 5/1-4x as power series.

OpenStudy (anonymous):

Let me write that up

OpenStudy (anonymous):

\[\frac{1}{1-x}=\sum _{n=0}^{\infty } x^n\] Replace x with 4x \[\frac{5}{1-4x}=\underset{n=0}{\overset{\infty }{5\sum }}(4x)^n\]

OpenStudy (anonymous):

hmm that's what i ended up with too, computer won't take it, must be a matter of syntax, thanks for the help!

OpenStudy (anonymous):

Yeah,I hate computer homework systems